Assume that when adults with smartphones are randomly​ selected,

4242​%

use them in meetings or classes. If

1414

adult smartphone users are randomly​ selected, find the probability that fewer than

44

of them use their smartphones in meetings or classes.

The probability is

nothing.

​(Type an integer or decimal rounded to four decimal places as​ needed.)
Answer #1

x~ Binomie 114, 0.42) 1x=y) - 1460 101423410-58) lo lo - 1 3 4 2
Answer #2

SOLUTION :


p = adults with smartphones = 42% = 0.42


q = adults without smartphones = 1 -  0.42 = 0.58


As per binomial distribution : For n = 14 :


P(x < 4) 


= P(0) + P(1) + P(2) + P(3) 

= ncr(14, 0)*(0.42)^0 *(0.58)^14 +ncr(14, 1)*(0.42)*(0.58)^13+ ncr(14, 2)*(0.42)^2 * (0.58)^12 + ncr(14, 3)*(0.42)^3 * (0.58)^11 


= 0.0961 (ANSWER).

answered by: Tulsiram Garg